OpenStudy (anonymous):

A model airplane can average 20mph with no wind for an 80 mile trip. Cycling into the wind, the bicyclist can cover 56 miles at the same time. Find the speed of the wind.

OpenStudy (anonymous):

model plane flies for 80/20 = 4 hours cyclist travels at 56/4 = 14 mph - thats all you can calculate with the information you have given

OpenStudy (anonymous):

you need to know the speed of the cyclist with no wind to find the wind speed

OpenStudy (anonymous):

that is all it says in the question/

OpenStudy (anonymous):

well there must be some info missing - maybe the cyclist's speed with no wind is the same as the airplane in which case speed of wind = 20-14 = 6 mph

OpenStudy (anonymous):

i think the cyclist's speed with no wind is the same as the airplane

OpenStudy (anonymous):

well if thats so 6 mph is your answer

OpenStudy (anonymous):

how did you get that then?